# Myopically Verifiable Probabilistic Certificates for Safe Control and Learning

### v1

@@ VII Experiments for Control
VII-A Algorithms for Comparison
We compare our proposed controller with three existing safe controllers designed for stochastic systems. Below, we present their techniques in the settings of this paper. We consider long-term safety in (LABEL:eq:long_term_safety) with fixed time horizon and time-invariant zero margin, i.e., $\mathbb{P}\left(\Phi_{X_{t}}(H)\geq 0\right)$.
  • Proposed controller: The safety condition is given by
    $D_{\mathbf{F}}(Z_{t},U_{t})\geq-\alpha(\mathbf{F}(Z_{t})-(1-\epsilon)),$ (100)
    where $\alpha>0$ is set to be a constant, and $\mathbf{F}$ is defined based on type 1 in (37).
  • Stochastic control barrier functions (StoCBF) [18]: The safety condition is given by
    $D_{\phi}(X_{t},U_{t})\geq-\eta\phi(X_{t}),$ (101)
    where $\eta>0$ is a constant. Here, the mapping $D_{\phi}:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}$ is defined as the infinitesimal generator of the stochastic process $X_{t}$ acting on the barrier function $\phi$, i.e.,
    $\begin{split}D_{\phi}(X_{t},U_{t}):=&A\phi(X_{t})\\ =&{\mathcal{L}}_{f}\phi(X_{t})+{\mathcal{L}}_{g}\phi(X_{t})U_{t}\\ &+\frac{1}{2}\text{tr}\left(\left[\sigma(X_{t})\right]\left[\sigma(X_{t})\right]^{\intercal}\operatorname{Hess}{\phi}(X_{t})\right).\end{split}$ (102)
    This condition constrains the average system state to move within the tangent cone of the safe set.
  • Probabilistic safety barrier certificates (PrSBC) [8]: The safety condition is given by
    $\mathbb{P}\left(D_{\phi}(X_{t},U_{t})+\eta\phi(X_{t})\geq 0\right)\geq 1-\epsilon,$ (103)
    where $\eta>0$ is a constant. This condition constrains the state to stay within the safe set in the infinitesimal future interval with high probability.
  • Conditional-value-at-risk barrier functions (CVaR) [58]: The safety condition is given by
    $\text{CVaR}_{\beta}\left(\phi(X_{t_{k+1}})\right)\geq\gamma\phi(X_{t_{k}})$ (104)
    where $\gamma\in(0,1)$ is a constant, $\{t_{0}=0,t_{1},t_{2},\cdots\}$ is a discrete sampled time of equal sampling intervals. This is a sufficient condition to ensure the value of $\text{CVaR}^{k}_{\beta}(\phi(X_{t_{k}}))$ conditioned on $X_{0}=x$ to be non-negative at all sampled time $t_{k\in\mathbb{Z}_{+}}$. The value of $\text{CVaR}^{k}_{\beta}(\phi(X_{t_{k}}))$ quantifies the evaluation made at time $t_{0}=0$ about the safety at time $t_{k}$.
VII-B Settings
@@
We run simulations with $\Delta t=0.1$ for all controllers unless otherwise specified. The initial state is set to $x_{0}=3$. For our controller, each Monte Carlo approximation uses $10000$ sampled trajectories. The parameters used are listed in Table I. Since the parameter $\alpha$ in the proposed controller has a similar effect as $\eta$ in StoCBF and PrSBC, we use the same values for these parameters in those controllers. The parameter $\epsilon$ is the tolerable probability of unsafe events both in the proposed controller and PrSBC, so we use the same values of $\epsilon$ for both algorithms for a fair comparison.
TABLE I: parameters used in simulation
Controller Parameters
Proposed controller $\alpha=1$, $\epsilon=0.1$, $H=10$
StoCBF $\eta=1$
PrSBC $\eta=1$, $\epsilon=0.1$
CVaR $\gamma=0.65$, $\beta=0.1$
VII-C Example Use Cases
VII-C1 System 1
@@
Fig. 1 shows the results in the worst-case setting. The proposed controller can keep the expected safe probability $\mathbb{E}[\mathbf{F}(X_{t})]$ close to $0.9$ all the time, while others fail to keep it at a high level with used parameters. A major cause of failure is due to the accumulation of rare event probability, leading to unsafe behaviors. This shows the power of having a provable performance for non-decreasing long-term safe probability over time. For comparable parameters, the safety improves from StoCBF to PrSBC to CVaR. This is also expected as constraining the expectation has little control of higher moments, and constraining the tail is not as strong as constraining the tail and the mean values of the tail.
@@ VII-C2 System 2
$f(x)=\begin{cases}2,\text{ if }x>1.5\\ -3,\text{ otherwise,}\end{cases}g(x)=\begin{cases}1,\text{ if }x>1.5\\ 0,\text{ otherwise.}\end{cases}$ (106)
This modification makes the system uncontrollable when the state $x$ reaches or is below $1.5$, even though $x=1.5$ remains safe. We keep noise magnitude $\sigma(X_{t})\equiv 2$, safe set ${\mathcal{C}}(0)=\left\{x\in\mathbb{R}:x-1\geq 0\right\}$, nominal controller $N(X_{t})=-2.5X_{t}$ the same. With the nominal controller, the system will reach $x=1.5$. Once the system reaches $x=1.5$, the autonomous system will be converging to the origin, which is unsafe. The switch in the dynamics and the nonlinear trap will get the system to unsafe regions even though the starting point is safe. This design will help illustrate the advantage of the predictive nature of our proposed controller compared to the existing myopic ones.
Fig. 3 shows the results in the worst-case setting. The proposed controller can keep the expected safe probability $\mathbb{E}[\mathbf{F}(X_{t})]$ close to $0.8$ all the time, while all other methods get 00 expected safe probability due to the nonlinear trap in the system. This shows the advantage of considering long-term safety instead of safety in the immediate next step.
@@
VII-D Neural Network-based Nominal Controller
We consider the linear control affine system (2) with $f(X_{t})\equiv A=2$, $g(X_{t})\equiv 1$, $\sigma(X_{t})\equiv 2$. The safe set is defined as in (105). We use a 2-layer neural network $h^{2}$ as the nominal controller, with $W_{1}=0.5$, $W_{2}=1$, $b_{1}=0.2$, $b_{2}=-0.5$. The activation function $\sigma_{\text{act}}$ is chosen to be the ReLU function. The nominal controller has the form
$N(X)$ $=\sigma_{\text{act}}\left(W_{2}\cdot\sigma_{\text{act}}(W_{1}X+b_{1})+b_{2}\right).$ (107)
In the safety probability estimation and safe control phase, we only have access to the value of $N(X)$ instead of the exact expression (107). We show the results for the proposed safe control method in Fig. 5. Since we can not write down the closed-loop dynamics of the system given the black-boxed neural network controller, none of the existing methods being compared in the previous section can be used. We can see that the nominal NN controller will yield unsafe behaviours, while with our proposed strategy long-term safety is ensured.
Figure 5: Results in the switching control setting with NN nominal controllers. (a) the average system state over 100 trajectories with its standard deviation. Red dotted line indicates the boundary of the safe set. (b) the empirical safe probability.
VIII Experiments for Reinforcement Learning
@@
where the probability is evaluated assuming $U_{t}\sim\pi(x_{t}),\forall t\in\{1,\cdots,H-1\}$.
VIII-A Policy Gradient
@@
We then apply the proposed safety filter to the system. We consider the safe set to be $\mathcal{C}=\{x\in\mathbb{R}:x-5>0\}$, i.e., we want to limit the state transition to any $x>5$ for safety concerns. As a result, we define the following safety filter which corresponds to the probabilistic safety certificate to the safest nominal policy $\bar{\pi}$:
@@
Here the safety filter outputs action ‘$-1$’ when $X>5$, otherwise outputs the nominal action $U$. This safety filter (88) can be acquired by running the proposed safe control strategy with $\bar{\pi}$ being the nominal controller. We conduct modified policy gradient described in (92) and the results are shown in Fig. 6 in orange lines. We can see that the learned policy will maximize the return, but will limit the state to safe regions thanks to the safety filter.
@@
VIII-B Q-learning
We consider the infinite time horizon problem so that Q function only depends on state and action. We set the discount function to be $\gamma=0.9$ so that the effective horizon is $10$. We set the initial state as $X_{0}=0$. Fig. 7 shows the change of Q function with and without the safety filter. We can see that the Q function will converge in 1500 iterations, and taking action $u=1$ will give the highest Q function value as expected. With the safety filter, the Q function converges to lower values, as the safety filter will limit the visitation of those high-rewarding but unsafe regions in the state space. If we consider greedy policy based on the learned Q function (take action that gives maximized Q function, i.e., $\pi^{Q}(x,u)=\arg\max_{u}Q(x,u)$), we should expect similar control behaviours as policy gradient with and without the safety filter.
@@ VIII-B Q-learning
IX Conclusion
This paper focuses on the problem of ensuring long-term safety with high probability in stochastic systems. The major challenge of this problem is the stringent tradeoffs between longer-term safety vs. computational burdens. To mitigate the tradeoffs, we explore how to impose forward invariance on a probability space. Even when set invariance on the state space is satisfied with high probability, long-term safety may not be guaranteed due to the accumulation of uncertainty and risk over time. In contrast, imposing probabilistic invariance on long-term probability allows myopic conditions/controllers to assure long-term design specifications. We then integrate this technique into both control and learning methods. The advantages of the proposed control and learning methods are demonstrated using numerical examples. While beyond the scope of this paper, probabilistic invariance generalizes to settings with unknown dynamics [24], multi-agent coordination [59], and latent risks with unobservable state [60]. The future direction of this research is to explore the probabilistic invariance technique for other control problems currently solved using a set-invariance-based approach, such as constrained optimal control, stability analysis, and robust control [1]. Given the similarity between set invariance and probabilistic invariance, we expect that probabilistic invariance techniques could provide new insights into these problems.

### v2

@@ 7 Experiments for Control
7.1 Settings
We compare our proposed controller with three existing safe controllers: stochastic control barrier functions (StoCBF) [20], probabilistic safety barrier certificates (PrSBC) [10], and conditional-value-at-risk barrier functions (CVaR) [77]. These baseline methods for comparison are widely acknowledged methods for safe control of stochastic systems, and have been used in safe trajectory synthesis and optimization [78], safe control of quadrotor swarms [79], risk-sensitive motion planning [80] and robot locomotion [81], etc. Note that all three methods for comparison assume the observation of the current state and the access to the system uncertainty model, which are the basic assumptions in our method as well.
@@ 7.1 Settings
We run simulations with $\Delta t=0.1$ for all controllers unless otherwise specified. The initial state is set to $x_{0}=3$. For our controller, each Monte Carlo approximation uses $10000$ sampled trajectories. Details on the methods for comparison and the used parameters can be found in Appendix .1 in the extended version of the paper submitted along with the supplementary materials.
7.2 Example Use Cases
7.2.1 System 1
@@
Fig. 1 shows the results in the worst-case setting. The proposed controller can keep the expected safe probability $\mathbb{E}[\mathbf{F}(X_{t})]$ close to $0.9$ all the time, while others fail to keep it at a high level with used parameters. A major cause of failure is the accumulation of rare event probability, leading to unsafe behaviors. This shows the power of having a provable performance for non-decreasing long-term safe probability over time. For comparable parameters, the safety improves from StoCBF to PrSBC to CVaR. This is also expected as constraining the expectation has little control of higher moments, and constraining the tail is not as strong as constraining the tail and the mean values of the tail.
@@ 7.2.2 System 2
[…]
This modification makes the system uncontrollable when the state $x$ reaches or is below $1.5$, even though $x=1.5$ remains safe. We keep noise magnitude $\sigma(X_{t})\equiv 2$, safe set ${\mathcal{C}}(0)=\left\{x\in\mathbb{R}:x-1\geq 0\right\}$, nominal controller $N(X_{t})=-2.5X_{t}$ the same. With the nominal controller, the system will reach $x=1.5$. Once the system reaches $x=1.5$, the state will to the origin, which is unsafe. The switch in the dynamics and the nonlinear trap will get the system to unsafe regions even though the starting point is safe. This design will help illustrate the advantage of the predictive nature of our proposed controller compared to the existing myopic ones.
Fig. 3 shows the results in the worst-case setting. The proposed controller can keep the expected safe probability $\mathbb{E}[\mathbf{F}(X_{t})]$ close to $0.8$ all the time, while all other methods get 0 expected safe probability due to the nonlinear trap in the system. This shows the advantage of considering long-term safety instead of safety in the immediate next step.
@@
Safe control results on neural network-based nominal controllers and 3-dimensional systems can be found in Appendix .3 and Appendix .4, respectively.
@@
8 Experiments for Reinforcement Learning
@@
where the probability is evaluated assuming $U_{t}\sim\pi(x_{t}),\forall t\in\{1,2,3,\cdots,H-1\}$.
8.1 Policy Gradient
@@
We then apply the proposed safety filter to the system. We consider the safe set to be $\mathcal{C}=\{x\in\mathbb{R}:x-5>0\}$, i.e., we want to limit the state transition to any $x>5$ for safety concerns. The safety filter corresponding to the safest nominal policy $\bar{\pi}$ is then given by
@@
Here the safety filter outputs action ‘1-1’ when $X>5$, otherwise outputs the nominal action $U$. This safety filter (71) can be acquired by running the proposed safe control strategy with $\bar{\pi}$ being the nominal controller. We conduct modified policy gradient described in (75) and the results are shown in Fig. 5 in orange lines. We can see that the learned policy will maximize the return, but will limit the state to safe regions thanks to the safety filter.
@@
8.2 Q-learning
We consider the infinite time horizon problem so that Q function only depends on state and action. We set the discount function to be $\gamma=0.9$ so that the effective horizon is $10$. We set the initial state as $X_{0}=0$ and number of episodes $N_{\text{eps}}=10$. Fig. 6 shows the change of Q function with and without the safety filter. We can see that the Q function will converge in 1500 iterations, and taking action $u=1$ will give the highest Q function value as expected. With the safety filter, the Q function converges to lower values, as the safety filter will limit the visitation of those high-rewarding but unsafe regions in the state space. If we consider greedy policy based on the learned Q function (take action that gives maximized Q function, i.e., $\pi^{Q}(x,u)=\arg\max_{u}Q(x,u)$), we should expect similar control behaviors as policy gradient with and without the safety filter.
@@ 8.2 Q-learning
9 Conclusion
This paper focuses on the problem of ensuring long-term safety with high probability in stochastic systems. The major challenge of this problem is the stringent tradeoffs between longer-term safety vs. computational burdens. To mitigate the tradeoffs, we explore how to impose forward invariance on a probability space. Even when set invariance on the state space is satisfied with high probability at each time, long-term safety may not be guaranteed due to the accumulation of uncertainty and risk over time. In contrast, imposing probabilistic invariance on long-term probability allows myopic conditions/controllers to assure long-term design specifications. We then integrate this technique into both control and learning methods. The advantages of the proposed control and learning methods are demonstrated using numerical examples. Beyond these contributions, probabilistic invariance naturally extends to a broad range of scenarios, including systems with unknown dynamics [31], multi-agent coordination [82], and latent risks with unobservable state [83].
Future research directions include generalizing the proposed approach to systems that are not affine control and to new control techniques. The requirement for affine control systems is common in many safe control techniques for deterministic systems. For stochastic systems, even for affine control systems, the safety conditions (e.g., the constraints for control barrier function conditions to hold with high probability) may not give rise to linear action constraints. In the proposed approach, the linearity of the safety constraint (26) requires affine control dynamics, while Theorem 1 generalizes to non-affine systems. This suggests the relaxation for affine control reduces to developing efficient techniques for evaluating the safety constraints (26) for non-affine-control systems. Additionally, the probabilistic invariance concept can be leveraged to tackle other control problems traditionally approached via set invariance, such as constrained optimal control, stability analysis, and robust control [1]. Since probabilistic invariance shares conceptual parallels with set invariance, it may provide new insights and alternative methodologies for such problems as well.